Hi can someone please help me solve for this, and or let me know if I'm on the right track? It is from the Financial Management: Principles and Applications text by Keown Chapter 5, question 5-1A:
(Compound interest) To what amount will the following investments accumulate?
a. $5,000 invested for 10 years at 10 percent compounded annually
Answer: 889,527.07
b. $8,000 invested for 7 years at 8 percent compounded annually
c. $775 invested for 12 years at 12 percent compounded annually
d. $21,000 invested for 5 years at 5 percent compounded annually

Is the answer I have for a.) correct? If not can you please explain what I did wrong and how I would do the other ones?
Thank you

Found 2 solutions by rfer, solver91311:
Answer by rfer(10417)

I am sending you A) so you can get started. I will look at the others.
Good rule of thumb is 10% interest doubles in seven years, so you can tell your answer is way off.
a)A=P*(1+r/n)^rt
A=5000(1.10)^10
A=5000(2.59374)
A=$12,968.71
------------------
b) A=8000(1+.08/1)^(1*7)
A=8000(1.08)^7
A=8000(1.713824269)
A=$13,710.59
--------------------
c) A=775(1+.12/1)^(1*12)
A=775(1.12)^12
A=775(3.895975993)
A=$3019.38
---------------------
d) A=21000(1+.05/1)^(1*5)
A=21000(1.05)^5
A=21000(1.276281563)
A=$26,801.19

Answer by solver91311(12126)



You need the following formula:



Where is the future value, is the beginning principal, is the annual interest rate (expressed as a decimal), and is the number of years.

So, for your first problem:



Start punching buttons on your calculator.

By the way, the answer you did give for #1 is WAY out in outer space somewhere. I tried but I can't figure out how you got there. Suffice it to say that to turn $5K into $890K in 10 years you would need to be earning something like 68% interest.

You can use the same formula for the other three problems.

The general formula for interest compounded at other periodicities is:



Where is the future value, is the beginning principal, is the annual interest rate, is the number of compounding periods per year and is the number of years.

Or, for continuous compounding:



Where is the future value, is the beginning principal, is the base of the natural logarithms (), is the annual interest rate, and is the number of years.


John